A ray of light is incident on an equilateral glass prism shows minimum deviation is 30°. Calculate the speed of light through prism?
1
Expert's answer
2019-08-19T09:49:40-0400

Let "A=60^{\circ}" be the angle of the prism, "D=30^{\circ}" be the angle of minimum deviation, then the refractive index "n" of the material of the prism can be calculated using the formula

"n=\frac{\sin (A+D)/2}{\sin A/2}=\sqrt{2}."

Let "c=300000\, km/h" be the speed of light in vacuum. The speed "v" of light through prism can be calculated using the formula

"v=c/n \approx 212132 \, km/h."